Write the equation of a line in standard form, passing through (-2,-3) and parallel to a line with the equation 3x-y=2
Answer by ankor@dixie-net.com(22740)

Write the equation of a line in standard form, passing through (-2,-3) and parallel to a line with the equation 3x-y=2
:
Since they are parallel both lines will have the same slope
Find the slope from the given equation, put it in the slope/intercept form
:
3x - y = 2
-y = -3x + 2
y = +3x - 2; y has to be positive, multiply by -1
the slope: m = 3
:
Find the line with coordinates -2,-3 and m=3;
use point slope formula y - y1 = m(x-x1)
:
y -(-3) = 3(x - (-2))
y + 3 = 3(x + 2)
y + 3 = 3x + 6
y = 3x + 6 - 3
y = 3x + 3
Standard form
-3x + y = 3
or
3x - y = -3